A fireman climbs a 9 m high ladder carrying
a 5.9999 in diameter hose that has a 1.2 in
diameter nozzle. The pump has an absolute
pressure of 4 atm .
What is the water velocity from the nozzle?
Assume that water is incompressible, its density
is 1000 kg/m3
, and 1 inch = 0.0254 m.
The acceleration of gravity is 9.8 m/s
2
and 1
atm = 1.01
Answer #1

Energy E = pressure P x volume V.

At the nozzle atm pressure also comes into play so we subtract atm pressure from absolute pressure.

P at nozzle = E/V = (4 - 1) x 101325 - gh x density = 215775 Pa or J/m3

Divide E/V by density to get E/m = 215775/1000 = 215.775 J/kg

v = sqrt(2E/m)

= sqrt(2 x 215.775) = 20.773 m/s